# When “Competency” in Reasoning Opens the Door to Vulnerability: Jailbreaking LLMs via Novel Complex Ciphers

### v2

@@ 5 Results and Discussion
Table 1: Decryption Success Rate (DSR) and Attack Success Rate (ASR) for baseline methods and our ACE (Automatic Cipher-based Encoding) technique across various LLMs. A ‘-’ in the DSR column indicates decoding the question was not required for that method. A ‘-’ in the ASR column signifies the technique applies only to open-source LLMs. The highest ASR for each model is highlighted in bold. The highest DSR is underlined.
@@
Table 2: Decryption Success Rate (DSR) and Attack Success Rate (ASR) for our LACE (Layered Automatic Cipher-based Encoding) method using Word Substitution Cipher + one encoding. The highest ASR for each model is highlighted in bold. The highest DSR is underlined.
In this section, we present the results of our jailbreaking methods as described in Section 4.
@@ 5.2 Performance of LACE
Table 2 shows the DSR and ASR of all the LLMs when evaluated against our second jailbreaking method, Layered Attacks using Custom Encryptions (LACE).
@@
Since GPT-4o is the only model able to successfully decode at a higher level, we see a significant increase in ASR for GPT-4o. Specifically, in cases of ROT-13 and Word Reversal Cipher, which are easier to solve, GPT-4o shows the highest ASR rate, up to 78%, along with 92% of DSR. It almost doubles the ASR of 40% from utilizing either one of Word Reversal Ciper and Word Substitution Ciper.This finding demonstrates our argument that bigger and more sophisticated models are more likely to be jailbroken while decrypting considerably complex ciphers.
For smaller models, although they weren’t able to decrypt the sentences properly, they may still decode enough to trigger a jailbreak attack. However, since their DSR is so low, the effectiveness of layering the ciphers is not observed on these models, suggesting the need to assign an appropriate level of cipher complexity based on the model’s capacity constraints.
@@ 6 Conclusion
In this study, we have identified an emerging vulnerability linked to the advanced reasoning capabilities of LLMs. While traditional jailbreak attacks often rely on single-layer encryption methods like Base64, we introduced two more sophisticated techniques: Attacks using Custom Encryptions (ACE) and Layered Attacks using Custom Encryptions (LACE). Within the ACE framework, we developed two novel ciphers—Grid Encoding and a Word-Substitution Cipher—that significantly increased the attack success rate (ASR) to 40% on GPT-4o. Additionally, with LACE, we enhanced the ASR by layering the Word-Substitution Cipher with other methods such as Word Reversal, resulting in a 38% improvement. These findings highlight the urgent need for more robust security measures in LLMs to anticipate and defend against increasingly complex attack vectors.
@@
While our methods are successful in jailbreaking LLMs and demonstrate the vulnerabilites that arise with the increasing capabilities of the model, it represents an early step in this area and has some limitations, including but not limited to the following:
@@
    AdvBench dataset is slightly biased towards the category of “Cyber Security”. Development of a new dataset that is more balanced across all 14 unsafe categories is left as a future work
@@
    Our work is limited to single-turn conversations, but in future, we would like to expand on multi-turn conversations.
Appendix A Encryption Ciphers
@@
Appendix B Priming Sentence
@@
Appendix C Harmful Taxonomy
@@
Appendix D GPT-4o-mini as a Judge

### v3

@@ 5 Results and Discussion
Table 1: DSR and ASR for baseline methods and our ACE technique across various LLMs on AdvBench. A ‘-’ in the DSR column indicates decoding the question was not required for that method. A ‘-’ in the ASR column signifies the technique applies only to open-source LLMs. The highest ASR for each model is highlighted in bold. The highest DSR is underlined.
@@
Table 2: DSR and ASR for our LACE method using Word Substitution Cipher + one encoding on AdvBench. The highest ASR for each model is highlighted in bold. The highest DSR is underlined.
@@ 5.2 Performance of LACE
Table 2 shows the DSR and ASR of all the LLMs when attacked using LACE.
@@
Since GPT-4o is the only model able to successfully decode at a higher level, we see a significant increase in ASR for GPT-4o. Specifically, for ROT-13 and Word Reversal Cipher, which are easier to solve, GPT-4o shows the highest ASR rate, up to 78%, along with 92% DSR. It almost doubles the ASR of 40% from using just one cipher. This finding demonstrates our argument that bigger and more sophisticated models are more likely to be jailbroken when decoding complex ciphers.
Smaller models on the other hand weren’t able to decrypt the sentences properly but they still decoded enough words of the query to understand the meaning and trigger an attack. However, since their DSR is so low, the effectiveness of layering the ciphers is not observed on these models, suggesting the need to assign an appropriate level of cipher complexity based on the model’s decoding capabilities.
@@ 6 Conclusion
In this study, we demonstrated that as LLMs become more adept at reasoning, they become more vulnerable to attacks. While traditional jailbreak attacks often rely on a single common cipher like Base64 or ROT-13, we build on these and introduced: Attacks using Custom Encryptions (ACE) and Layered Attacks using Custom Encryptions (LACE). Within the ACE framework, we developed two novel ciphers—Grid Encoding and a Word-Substitution Cipher—that significantly increased the attack success rate (ASR) to 40% on GPT-4o. Additionally, with LACE, we enhanced the ASR by layering the Word-Substitution Cipher with other methods such as Word Reversal, resulting in a 38% improvement. These findings highlight the urgent need for more robust security measures in LLMs to anticipate and defend against increasingly complex attack vectors.
@@
While our methods are successful in jailbreaking LLMs and demonstrate the vulnerabilites that arise with the increasing reasoning capabilities of the model, it represents an early step in this area and has some limitations, including but not limited to the following:
@@
    We use GPT-4o-mini to evaluate whether the response is “Safe” or “Unsafe”. This makes our evaluation imperfect, as it misclassifies some responses. More discussion is present in Appendix E.
@@
    Our work is limited to single-turn conversations, but in the future, we would like to expand on multi-turn conversations.
Appendix A Results on JailbreakBench
DSR ASR DSR ASR DSR ASR DSR ASR
Priming Sentence - 0.00 - 0.01 - 0.46 - 0.48
PAIR - 0.06 - 0.17 - 0.16 - 0.12
AutoDAN - - - - - 0.80 - 0.89
ArtPrompt 0.01 0.43 0.01 0.37 0.00 0.26 0.00 0.40
LeetSpeak 0.85 0.49 0.86 0.37 0.79 0.64 0.04 0.59
Base64 0.96 0.51 0.81 0.44 0.81 0.55 0.00 0.23
ROT-13 0.81 0.52 0.00 0.15 0.00 0.48 0.00 0.21
Keyboard Cipher 0.09 0.38 0.00 0.26 0.00 0.10 0.00 0.15
Upside Down 0.65 0.52 0.00 0.31 0.00 0.23 0.00 0.21
Word Reversal 0.87 0.56 0.03 0.39 0.08 0.46 0.01 0.42
Grid Encoding 0.36 0.59 0.00 0.17 0.00 0.28 0.00 0.08
Word Substitution 0.98 0.77 0.86 0.51 0.86 0.62 0.78 0.79
Table 3: DSR and ASR for baseline methods and our ACE technique across various LLMs on JailbreakBench. A ‘-’ in the DSR column indicates decoding the question was not required for that method. A ‘-’ in the ASR column signifies the technique applies only to open-source LLMs. The highest ASR for each model is highlighted in bold. The highest DSR is underlined.
Attack Method GPT-4o Gemini-1.5-Flash Llama-3.1-70B Llama-3.1-8B
Word Substitution Cipher + DSR ASR DSR ASR DSR ASR DSR ASR
ROT-13 Sentence 0.77 0.74 0.01 0.40 0.04 0.47 0.00 0.54
Mappings 0.34 0.71 0.00 0.33 0.00 0.35 0.00 0.28
Both 0.26 0.64 0.00 0.13 0.00 0.27 0.00 0.22
Keyboard Sentence 0.10 0.50 0.00 0.36 0.00 0.30 0.00 0.40
Mappings 0.00 0.27 0.00 0.34 0.00 0.17 0.00 0.24
Both 0.00 0.17 0.00 0.16 0.00 0.14 0.00 0.17
Grid Encoding Sentence 0.30 0.72 0.01 0.44 0.01 0.52 0.00 0.30
Mappings 0.06 0.37 0.01 0.22 0.00 0.20 0.00 0.08
Both 0.01 0.17 0.00 0.13 0.00 0.13 0.00 0.07
Word Reversal Sentence 0.90 0.74 0.17 0.48 0.26 0.61 0.04 0.62
Mappings 0.62 0.69 0.04 0.43 0.10 0.44 0.00 0.38
Both 0.70 0.69 0.01 0.28 0.02 0.22 0.00 0.23
Table 4: DSR and ASR for our LACE method using Word Substitution Cipher + one encoding on JailbreakBench. The highest ASR for each model is highlighted in bold. The highest DSR is underlined.
Table 3 and 4 show the results of ACE and LACE respectively on JailbreakBench (Chao et al., 2024). Similar to results discussed in Section 5, we see that using novel user-created ciphers are more effective in jailbreaking LLMs compared to ciphers that are widely studied in the literature.
Appendix B Encryption Ciphers
@@
Appendix C Priming Sentence
@@
Appendix D Harmful Taxonomy
@@
Appendix E GPT-4o-mini as a Judge
@@
To verify whether GPT-4o-mini is capable of classifying properly, we randomly sample 50 judgments and ask three annotators to verify it. The annotators are graduate students in Computer Science and are proficient in English. They are volunteers who are not a part of our work and were told beforehand that no monetary compensation would be provided. Table 5 provides the inter-annotator agreement between the annotators and GPT-4o-mini.
Ann. 1 Ann. 2 Ann. 3 Judge
Ann. 1 - 0.84 0.84 0.68
Ann. 2 0.84 - 0.76 0.76
Ann. 3 0.84 0.76 - 0.60
Judge 0.68 0.76 0.60 -
Table 5: Cohen-Kappa score between the three annotators and Judge (GPT-4o-mini) for classifying responses as “Safe” or “Unsafe”.
We observe an average Cohen-Kappa score of 0.68 which signifies Substantial Overlap. As noted in the Limitations Section, GPT-4o-mini is not a perfect judge, but we believe it mimics an actual human well in classifying responses as “Safe” or “Unsafe”.